OpenStudy (anonymous):

David's company reimburses his expenses on food, lodging, and conveyance during business trips. The company pays $60 a day for food and lodging and $0.65 for each mile traveled. David drove 600 miles and was reimbursed $3,390. Part A: Create an equation that will determine the number of days x on the trip. (3 points) Part B: Solve this equation justifying each step with an algebraic property of equality. (6 points) Part C: How many days did David spend on this trip? (1 point)

OpenStudy (danjs):

Total Back = Food + Lodging + traveling distance

OpenStudy (anonymous):

$3390 = 0.65*600 + 60*x ?

OpenStudy (danjs):

looks good

OpenStudy (anonymous):

for part a?

OpenStudy (anonymous):

alrght

OpenStudy (anonymous):

what about part b?

OpenStudy (danjs):

Looks like you have to explain every step you do to isolate x on one side

OpenStudy (danjs):

these look familiar, i think they want those words http://www.mathwords.com/e/equation_rules.htm

OpenStudy (anonymous):

would that be good to submit?

OpenStudy (danjs):

umm..

OpenStudy (danjs):

flip equation, make a not of the symmetric property a = b then b = a

OpenStudy (danjs):

**note

OpenStudy (anonymous):

So what would i say?

OpenStudy (danjs):

I would not use subtraction or division, they are just a fancy word for addition and multiplication

OpenStudy (danjs):

--'Subtract 390 from both sides' i would put 'ADDITION Property of (-6) to both sides'

OpenStudy (danjs):

no need to put simplified , you just took care of the non alebra numbers

OpenStudy (anonymous):

ok, now its good to submit?

OpenStudy (danjs):

---Instead of divide by 60 I would put ; Multiplication property of Equality" multiplied by 1/60

OpenStudy (anonymous):

that it??

OpenStudy (danjs):

i think so

OpenStudy (danjs):

less you want to state each property after you name it, like IF a = b, then c*a = c*b , where c=(1/60) for the multiplication property step